You have predicted the following returns for stocks A and B in three possible states of nature. What is expected return for stock A?

State
Probability
A
B
Boom
0.1
0.20
0.30
Normal
0.4
0.10
0.20
Recesion
0.5
0.05
0.07

Select one:

a. 8.5%

b. 10.5%

c. 8.7%

d. 9.5%
Answer #1
Boom Normal Recession
State Probability 0.10 0.40 0.50
A 0.20 0.10 0.05
B 0.30 0.20 0.07

Expected return for stock A = (p1 * r1) + (p2 * r2) + ................ + (pn * rn)
Where p implies Probability of that event happening and r implies Return the stock generates when that event happens.

Expected Return = (0.10 X 0.20) + (0.40 X 0.10) + (0.50 X 0.05)
  = 0.02 + 0.04 + 0.025 = 0.085 or 8.5%

Therefore, the correct option is A (8.5%)